OpenStudy (anonymous):

Which of the following is the radical expression of a^(5/7)

jimthompson5910 (jim_thompson5910):

Use the idea that \[\Large x^{\frac{m}{n}} = \sqrt[n]{x^m}\] to get \[\Large a^{\frac{5}{7}} = \sqrt[7]{a^5}\]

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome